2Pacfan10:

6 stamps cost 30¢.How many stamps can be bought for 45¢

Angle:

If 6 stamps costs 30 cents how much does one stamp cost?

2Pacfan10:

idk

Angle:

What number times 6 equals 30?

aevans182:

5

Angle:

cool, so each stamp costs 5 cents

Angle:

If each stamp costs 5 cents, how many can you buy with 45 cents?

aevans182:

Can I say it?

2Pacfan10:

okie ty

snowflake0531:

Write a proportion

aevans182:

It's 5+4

snowflake0531:

6 stamps is 30, so 6/30, x stamps is 45, so x/45

snowflake0531:

so 6/30 = x/45

snowflake0531:

but you can simplify the first to 1/5 = x/45

snowflake0531:

so then, cross multiply 5x=45

Angle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @aevans182 It's 5+4 \(\color{#0cbb34}{\text{End of Quote}}\) Sure that does get them to the correct numerical answer, but it does not show them the correct ideas to use in order to get the right answer by themselves in the future

snowflake0531:

and since i can't give a direct answer, 5x=45 solve for x xD

snowflake0531:

thx katrin for the medal

aevans182:

Ok. I thought it would be ok bc you already explained it @Angle

snowflake0531:

oop

2Pacfan10:

so its 9?

Angle:

yes :)

2Pacfan10:

yay :)

snowflake0531:

lol